OpenStudy (ao77ao77):

Please Help y = 2x + 4 2x - y = 6 Solve the system of equations using substitution. A) (2,8) B) (2,-8) C) (10,24) D) no solution

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge y=2x+4$$ $$\Huge 2x–y=6$$ $$\huge 2x –(2x+4)=6$$$

OpenStudy (vheah):

That's correct, and then you distribute the negative inside the parentheses giving you 2x-2x-4=6 If you try to solve for x you will find 2x-2x =0 leaving yu with -4=6 but -4 is not equal to 6, therefore there is no solution.

OpenStudy (vheah):

another thing you could do, although it might take a while, is plug in each point to the system and see if you get an equal answer for both. If not, then it's D.

OpenStudy (skullpatrol):

$$\Huge –(a+b)=(–a)+(–b)$$